OpenStudy (zuha):

Use the complex conjugate to find the absolute value of 8+12i.

OpenStudy (zuha):

I'm not completely sure what a complex conjugate is to begin with.

OpenStudy (518nad):

basically u just take the same complex number but u make the i part negative

OpenStudy (518nad):

(8+12i)*=8-12i

OpenStudy (zuha):

So basically you cancel it out?

OpenStudy (518nad):

zz*=|z|^2 z=(8+12i) (8+12i)(8+12i)* =8^2+12^2=64+144=208

OpenStudy (zuha):

Where did the z come from?

OpenStudy (518nad):

|z|^2=208 |z|=sqrt208

OpenStudy (518nad):

z represents a complex number

OpenStudy (518nad):

(a+ib ) x (a-ib) =? see what happens when u multiply these numbers together

OpenStudy (518nad):

a+ib represents a a complex number z and a-ib represents the complex conjugate of z

OpenStudy (518nad):

all complex numbers are of the form a+ib

OpenStudy (zuha):

Is there a purpose for the asterisk here "(8+12i)(8+12i)*"?

OpenStudy (518nad):

the asterisk means complex conjugate

OpenStudy (518nad):

(8+12i)(8+12i)* = (8+12i)(8-12i)

OpenStudy (518nad):

use your distributive law of multiplication to expand those brackets out

OpenStudy (zuha):

When we multiply them together we get a^2-ib^2 I believe?

OpenStudy (zuha):

Alright, that makes sense, just read through everything you said, thank you for your help :)